# Exponents and front fluctuations in the quenched Kardar-Parisi-Zhang universality class of one and two dimensional interfaces

### v1

@@ VII Summary and conclusions
  • We have computed the critical force for depinning $F_{c}$ and the velocity exponent $\theta$, as well as the scaling exponent $\delta$ which rules the dependence of the front velocity with time at $F=F_{c}$.
@@
  • From the computed exponents, we have verified the scaling relations $\beta+\delta=1$ and $\beta z=\alpha$. In this respect, we remark that the exponents $\alpha$ and $z$ were obtained directly from the data, without resorting to scaling relations.
@@
  • Finally, we have assessed the occurrence of asymptotically time and system-size independent PDF for the front fluctuations in the growth regime. We have described the shape of these distributions in terms of their skewness, excess kurtosis, and their tail exponents. Specifically, the tails of these PDFs were found to differ noticeably from those commonly found in the disorder-dominated phases of other random systems.
Overall, the results reported herein show that the automaton versions reasonably well the critical behavior of the qKPZ equation without exhibiting convergence problems or instabilities of the integration schemes.
@@ Appendix B Correlation function
Figure 10 shows two sample unscaled $C_{2}(\boldsymbol{r},t)$ functions, arbitrarily taken as representatives for both the 1D (a) and 2D (b) models, for several times, and fixed $L$. As previously mentioned, no signs of anomalous scaling (typically, a time-dependent shift of the $C_{2}(\boldsymbol{r},t)$ vs. $r$ curves for increasing $t$, see e.g. Ref. [Barreales2023]) are observed, which is a sound demonstration of the validity of the FV scaling for both dimensionalities.

### v2

@@ VII Summary and conclusions
  • We have computed the critical force for depinning $F_{c}$ and the velocity exponent $\theta$, as well as the scaling exponent $\delta$ which rules the dependence of the front velocity with time at $F=F_{c}$. To our knowledge, the value of $F_{c}$ for 2D is reported here for the first time.
@@
  • From the computed exponents, we have verified the scaling relations $\beta+\delta=1$ and $\beta z=\alpha$. In this respect, we remark that the exponents $\alpha$ and $z$ were obtained directly from the data, without resorting to scaling relations. Actually, to the best of our knowledge, this is the first time that the entire set of critical exponents has been computed directly from simulations in this system for both dimensionalities.
@@
  • Finally, we have assessed the occurrence of asymptotically time and system-size independent PDF for the front fluctuations in the growth regime. We have described the shape of these distributions in terms of their skewness, excess kurtosis, and their tail exponents. Specifically, the tails of these PDFs were found to differ noticeably from those commonly found in the disorder-dominated phases of other random systems. Again, this study has been performed for the first time both in 1D and 2D, to our knowledge.
Overall, the results reported herein show that the automaton versions describe reasonably well the critical behavior of the qKPZ equation without exhibiting convergence problems or instabilities of the integration schemes.
@@ Appendix B Correlation function
Figure 10 shows two sample unscaled $C_{2}(\boldsymbol{r},t)$ functions, arbitrarily taken as representatives for both the 1D (a) and 2D (b) models, for several times, and fixed $L$. As previously mentioned, no signs of anomalous scaling (typically, a time-dependent shift of the $C_{2}(\boldsymbol{r},t)$ vs. $r$ curves for increasing $t$, see e.g. Ref. [50]) are observed, which is a sound demonstration of the validity of the FV scaling for both dimensionalities.